It is presentation day in class and your instructor is drawing names from a hat to determine the order of the presentations. if there are 18 students in the class, what is the probability that the first 2 presentations will be by Kelly and harmony, in that order? express your answer as a fraction in lowest terms or a decimal rounded to the nearest millionth.
1
Expert's answer
2020-10-21T16:49:27-0400

Given : n= 18

r= 2

solution:

since order matters here, we use permutation.

formula:

npr= "\frac{n!}{(n-r)!)}"

Hence 18P2 ="\frac{18!}{(18-2)!}" = 306 ways in total

number of ways selecting two = 1(since order matters)


Hence required probability = "\frac{1}{306}"